Grade 12Analytical Geometry

please answer fast with solution
the length of latus rectum of parabola 2y^2+3y+4x-2 is???

Mrunal Sonawane
7 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
It can be written as 
 
2 (y^2 + 3/2y) = 2 – 4x
 
y^2 + 3/2 y + 9/16 = 1 – 2x + 9/16
 
(y + 3/4)^2 = – 2( x – 25/16)
 
hence length of latus 4ectum = 2